given an interest rate of 10 percent, compute the year 7 future value if deposits of $2800 and $3800 are made in years 2 and 3 respectively and a withdrawal of $925 is made is year

5
Answer #1

Year 7 future value is $ 8,953.76

Working:

Calculation of year 7 value of cash flows:
Year Beginning cash flow Interest Addition in cash flow Ending cash flow
a b=a*10% c d=a+b+c
1 0 0 0 0
2 0 0 $ 2,800.00 $ 2,800.00
3 $ 2,800.00 $ 280.00 $ 3,800.00 $ 6,880.00
4 $ 6,880.00 $ 688.00 0 $ 7,568.00
5 $ 7,568.00 $ 756.80 $   -925.00 $ 7,399.80
6 $ 7,399.80 $ 739.98 0 $ 8,139.78
7 $ 8,139.78 $ 813.98 0 $ 8,953.76